# Exploring filament galaxies using AstroSat/ UVIT

### v1

@@ 4 Properties of filament galaxies
We study the star-forming, nuclear, and structural properties of all 58 galaxies in the four filaments. As shown in Figure 2, only a portion of these filaments were observed with UVIT. Therefore, we create two subsets among the total galaxies observed with UVIT (within 1 Mpc from the nearest filament axis): the FUV detected and FUV undetected galaxies. The histogram shown in the upper left panel of Figure 3 highlights the cross-section distribution of galaxies inside filaments from the filament axis for the entire sample and both subsets. The filaments seem to be more dense in the inner region. Interestingly, we detect no galaxy in our FUV observation up to D 0.4 Mpc.
@@ 4.1 Color-magnitude diagram
We use optical gr vs. M color-magnitude diagram (CMD) to classify all 58 filament galaxies into the blue cloud, red sequence, and green valley galaxies (Blanton et al., 2005; Salim, 2014). The blue cloud contains star-forming galaxies, mostly disk-dominated spirals, and irregulars. The red sequence comprises star formation quenched galaxies. Lenticular and elliptical galaxies mostly dominate this region. The intermediate region between the two modes of distribution is known as the green valley, where galaxies with intermediate star formation histories lie. The parallel lines dividing the CMD are given by Blanton et al. (2005). The background galaxies in the sample are taken from SDSS DR 16 within 0.1.
As seen in the upper right panel of Figure 3, our filament sample is spread across the blue cloud and green valley region. We detect a comparable population of star-forming galaxies and green valley galaxies in our FUV survey. On the contrary, we only detected star-forming and blue galaxies in our previous study of the Bootes void (Pandey et al., 2021). We highlight the dominance of green valley galaxies in filaments which contain a mix of stellar population. Such galaxies are in transition phase where the star formation is declining but not entirely ceased. Galaxy interactions and mergers, feedback from active-galactic nuclei (AGN) are few physical mechanisms that can drive a galaxy to this state (Salim, 2014). The galaxies present in filaments studied by us are brighter than M 20 mag. However, previous reports have confirmed the presence of faint dwarf galaxies in filament environment as well (Lee et al., 2021).
@@ 4.2 BPT diagram
The BPT (Baldwin, Phillips, and Terlevich, Baldwin et al., 1981) diagram is ubiquitously used to classify and understand the ionization sources of galaxies based on their emission line ratios. It provides insights into the dominant mechanisms that power the emission lines observed in the spectra of galaxies. We use the line flux ratio of [O III]/H vs. [N II]/ H. The diagram is divided into the three regions containing purely star-forming or ‘HII region’ galaxies, composite galaxies, and galaxies with AGN based on prescription given by Kewley et al. (2001) and Kauffmann et al. (2003a).
The emission line fluxes used in the diagram were corrected for internal extinction prior to use. The reddening value was determined using H and H emission lines following the Balmar decrement method given by Osterbrock (1989). It can be seen in the bottom left panel of Figure 3 that filaments contain a mix of populations in terms of nuclear activity. All purely star-forming galaxies in the FOVs were detected in FUV observation. We also detect a few galaxies with Active Galactic Nuclei (AGN) in our UV survey. The AGN-host filament galaxies are star-marked in Table 2.
@@ 4.3 Morphological classification using non-parametric methods
here, X represents all pixel value , is the average of all X, is total number of pixels. All pixel values are arranged in an increasing order before calculating the . For a homogeneous system, converges to 0, whereas = 1 when the entire light of a galaxy is concentrated in a pixel (Glasser, 1962).
The parameter corresponds to the second-order moment of light for the brightest region in the galaxy (Lotz et al., 2004). The parameter is calculated by summing over the brightest pixel values () until the sum corresponds to 20% of the total galaxy flux. This value is normalized by M corresponding to the second-order moment of the total light in a galaxy. These quantities are defined as:
here, x, y corresponds to galaxy central coordinates.
@@ 5 FUV emission in filament galaxies
The FUV SFRs would serve as an excellent tool to identify the most recently star-forming galaxy in our sample. As highlighted in the previous section, apart from UVIT-identified galaxies, we find a set of filament galaxies that do not emit in FUV above a threshold of 1.5. These galaxies are termed as FUV undetected. We perform fixed aperture photometry with an aperture radius of 1.5 on the FUV images at the centroid positions of the undetected galaxies. The magnitudes, thus obtained, were aperture corrected up to the saturation radius of a point source. The average aperture corrected magnitudes were approximately 1 - 1.5 mag higher than the FUV 3 limiting magnitude. These magnitudes are used in the subsequent analysis in this work.
@@
We calculated internal extinction corrected FUV SFRs for all (UVIT detected and undetected) filament galaxies using following relation taken from Kennicutt (1998):
@@
Where Lis the intrinsic FUV luminosity. We determine the value of internal extinction by the method discussed earlier in section 4.2. The ratio between stellar-to-gas attenuation is assumed to be 0.44. The median FUV SFRs of UVIT detected filament galaxies is 1.0 M yr, similar to the Milky Way galaxy. The stellar mass of filament galaxies were procured from GALEX-SDSS-WISE Legacy Catalog (GSWLC, Salim et al., 2016). Table 2 contains the FUV magnitudes, SFRs, stellar mass, and D for all filament galaxies detected in UVIT. Figure 4 shows FUV specific star formation rates (sSFR ) vs. M distribution for all filament galaxies targeted in our UVIT survey. We use aperture-corrected FUV magnitudes for calculating FUV SFRs in the case of UVIT undetected galaxies. The color bar indicates the D for each galaxy.
@@
As seen in the figure, we find two distinct populations of galaxies in the filaments - star-forming and quiescent. A majority of the UVIT undetected galaxies lie closer to the filament spine. The non-detection of galaxies near the spine could be associated with the extra-galactic gas supply cutoff deep inside filaments suggested by Kuutma et al. (2017), which may lead to a dearth of recent star formation in the region. We find a weak positive trend between FUV sSFR and D. A similar trend of decreasing optical color index is reported in the literature (Kuutma et al., 2017; Singh et al., 2020). The relation is extensively discussed in the following section.
@@
The massive filament galaxies (M 10 M) detected in our FUV survey are mostly star-forming. Some of these galaxies also host an AGN. However, it is well-established that several stellar mass quenching processes, viz. gravitation quenching (Genzel et al., 2014), AGN feedback (Croton et al., 2006), active in massive galaxies lead to the cessation of the ongoing star formation. Intriguingly, a few studies on filaments establish that the inter-filamentary gas can replenish the HI content of filament galaxies (Kereš et al., 2005). Kleiner et al. (2017) reported the evidence of cold-mode gas accretion in several galaxies. This effect is more prominently observed in high-mass galaxies due to their deep gravitational potential well. As detected in our study, the accumulated cold gas can further ignite the star formation activities in the massive filament galaxies. However, the likelihood of gas accretion in the sampled filament galaxies needs further investigation.
5.1 Filament vs. clusters galaxies
One of the key objectives of this work is to investigate whether the star-forming properties of filament galaxies are different from those existing in other large-scale environments. Most of the galaxies detected in our FUV survey lie on the massive side of the stellar mass range. We do not include void galaxies in this analysis as voids are mostly dominated with low-stellar mass galaxies (Kreckel et al., 2012; Pandey et al., 2021). Galaxy clusters are another LSS known to host the most massive galaxies in the Universe. We compare two aspects of the FUV emission from cluster and filament galaxies. Firstly, we are interested in studying the gradient in FUV sSFRs of galaxies inside filaments and clusters as a function of their distances from the centers of the respective structures. Moreover, we are motivated to compare the FUV emission in massive filament and cluster galaxies.
@@
We use the RedMaPPer catalog (Rykoff et al., 2016) for identifying cluster galaxies in the nearby Universe (z 0.1). We select galaxies with cluster membership probability greater than 50%. The FUV magnitudes of cluster members were obtained from the GALEX AIS catalog. The galaxies with FUV magnitude error less than 0.3 mag are included in the sample. Using the FUV magnitudes, we calculate internal dust extinction FUV SFR using Equation 4. The value of internal reddening in the galaxies is deduced using the Balmar decrement ratio method (see Section 4.2). The stellar masses corresponding to the galaxies were taken from GSWLC. Forth applying all the cuts, we obtain a small sample of 33 cluster galaxies between a mass range of 10 M to 10 M.
Figure 5 depicts the dependence of FUV sSFR on the distance of filament spine and cluster center for filament and cluster galaxies, respectively. We include FUV undetected galaxies in the analysis for a coherent interpretation. It can be clearly seen that the recent star formation for filament galaxies decreases on approaching the filament spine. On the other hand, cluster galaxies show little-to-no dependence between FUV sSFRs and distance from the cluster center. Previously done studies pointed out that for nearby galaxies (z 0), the SFR and sSFR do not show any significant variation with clustercentric distance (Laganá & Ulmer, 2017). Even after including FUV undetected filament galaxies in the study, the correlation between sSFR and D remains unaltered. The results highlight the difference in evolution galaxies depending on the large-scale environment.
Figure 5 qualitatively shows that the average FUV sSFR of filament galaxies detected in FUV observation is greater than the cluster galaxies. We perform the Kolmogorov–Smirnov (KS) test and the Anderson-Darling (AD) test on the distribution of FUV sSFR for cluster and filament galaxies in a fixed stellar mass bin. We only selected galaxies having stellar mass greater than 10 M. The p-values corresponding to both tests were much less than 0.05, whereas KS-statistic (= 0.73) and AD-statistic (= 7.2) were high on comparing the FUV sSFR of UVIT detected filament and cluster galaxies. We combine the aperture-corrected FUV sSFR of undetected filament galaxies with the detected galaxies and perform the same statistical tests. The p-value increases to 0.02, less than the significance level of 0.05. Hence, we conclude that the FUV sSFR of massive filament and cluster galaxies belong to different parent populations. The statistical tests imply that the star formation process is more rapid in filaments than clusters. The observed results could arise from a slower gas depletion time scale in the former environment than the latter one, as pointed out by Castignani et al. (2022). Alternatively, the cold mode gas accretion from inter-filamentary channels or gas-rich mergers could rejuvenate the star formation in massive galaxies.
@@ 5.2 FUV morphology of filament galaxies
The FUV morphology of galaxies is an efficient probe to identify the recently star-forming regions in them. We utilize our well-resolved UVIT observation to investigate the FUV morphology of filament galaxies and compare them with their optical counterparts. Figure 6 illustrates FUV images of all UVIT-detected filament galaxies. The blue contours on the figure show the SDSS-r band profile of galaxies. The three contour levels in blue represent 10, 5, and 1.5 level, where represents the mean standard deviation in the SDSS r-band observation. The surface brightness of the outermost contour is 24.14 mag arcsec. The galaxies are serially assigned a number () according to Table 2.
@@
We show FUV and optical centroids of the galaxies using cyan and red crosses, respectively. The young stellar population inside galaxies represented by FUV emission mostly aligns with the main stellar population traced by optical emission. On the contrary, a similar analysis done for cluster galaxies by Mahajan et al. (2022) highlights that the FUV morphology for a sample of cluster galaxies was skewed relative to their optical emission primarily due to ram-pressure stripping (Gunn & Gott, 1972) rampantly observed in the cluster environment. The result establishes that filament galaxies may not witness extreme environment disturbances, which are prevalent in clusters. However, the FUV emission appears clumpy in the filament galaxies and shows considerable size variation. For a few galaxies, the emission is asymmetric within the galaxy, e.g., . Moreover, we observe FUV emission in some galaxies beyond the outermost optical contour, e.g., . These observations hint at the possibility of recent gas accretion episodes. A thorough comparison between deep optical and UV observation of the filament galaxies is required for confirming the existence of extended-UV disks (Thilker et al., 2007).
@@
As discussed in the previous section, massive galaxies in filament exhibit a faster rate of recent star formation than those in the clusters. The FUV disk sizes of most of the massive filament galaxies (M 10 M) in the sample are slightly extended or comparable to their optical emission. These galaxies are mostly detected in the outer edge of filaments (D Mpc). Previous studies establish that high stellar mass galaxies primarily reside closer to the filament spine (Chen et al., 2017). In our survey, massive galaxies near the filament axis were undetected or detected with weak UV emission (see in Figure 6; also see upper left panel of Figure 3). Some moderate stellar mass galaxies ( 10 M) also show extended UV emission in our sample (e.g., 11). Previous discussion in this work hints towards a gradient in the level of recent star formation as we progress from the filament axis to its outer edges. An earlier study analyzing the HI gas content of galaxies in the filament environment reported that at fixed stellar mass and color, the galaxies closer to the filament spine are increasingly HI deficient (Odekon et al., 2018). This finding corroborates our result wherein the galaxies closer to the spine do not show a signature of recent star formation.

### v2

@@ 4 Properties of filament galaxies
We study the star-forming, nuclear, and structural properties of all 58 galaxies in the four filaments. As shown in Figure 2, only a portion of these filaments were observed with UVIT. Therefore, we create two subsets among the total galaxies observed with UVIT (within 1 Mpc from the nearest filament axis): the FUV detected and FUV undetected galaxies. The histogram shown in the upper left panel of Figure 3 highlights the cross-sectional distribution of galaxies inside filaments from the filament axis for the entire sample and both subsets. The filaments seem to be more dense in the inner region. We emphasize that we detect no galaxy in our FUV observation up to Dfil 0.4 Mpc .
@@ 4.1 Color-magnitude diagram
We use optical gr vs. Mr color-magnitude diagram (CMD) to classify all 58 filament galaxies into the blue cloud, red sequence, and green valley galaxies (Blanton et al., 2005; Salim, 2014). The blue cloud contains star-forming galaxies, mostly disk-dominated spirals, and irregulars. The red sequence comprises star formation quenched galaxies. Lenticular and elliptical galaxies mostly dominate this region. The intermediate region between the two modes of distribution is known as the green valley, where galaxies with intermediate star formation histories lie. The parallel lines dividing the CMD are given by Blanton et al. (2005). The background galaxies in the sample are taken from SDSS DR 16 within 0.1.
As seen in the upper right panel of Figure 3, our filament sample is spread across the blue cloud and green valley region. We detect a comparable population of star-forming galaxies and green valley galaxies in our FUV survey. On the contrary, we only detected star-forming and blue galaxies in our previous study of the Bootes void (Pandey et al., 2021, 2022). We highlight the dominance of green valley galaxies in filaments that contain a mix of stellar populations. Such galaxies are in a transition phase where the star formation is declining but not entirely ceased. Galaxy interactions and mergers, feedback from active galactic nuclei (AGN) are a few physical mechanisms that can drive a galaxy to this state (Salim, 2014). The galaxies present in filaments studied by us are brighter than Mr 20 mag. However, previous reports have confirmed the presence of faint dwarf galaxies in filament environments as well (Lee et al., 2021).
@@ 4.2 BPT diagram
The BPT (Baldwin, Phillips, and Terlevich, Baldwin et al., 1981) diagram is ubiquitously used to classify and understand the ionization sources of galaxies based on their emission line ratios. It provides insights into the dominant mechanisms that power the emission lines observed in the spectra of galaxies. We use the line flux ratio of [O III]/Hβ vs. [N II]/ Hα. The diagram is divided into the three regions containing purely star-forming or ‘HII region’ galaxies, composite galaxies, and galaxies with AGN based on prescription given by Kewley et al. (2001) and Kauffmann et al. (2003a).
The emission line fluxes used in the diagram were corrected for internal extinction before use. The reddening value was determined using Hα and Hβ emission lines following the Balmar decrement method given by Osterbrock (1989). It can be seen in the bottom left panel of Figure 3 that filaments contain a mix of populations in terms of nuclear activity. All purely star-forming galaxies in the FOVs were detected in FUV observation. We also detect a few galaxies with AGN in our UV survey. The AGN-host filament galaxies are star-marked in Table 2.
@@ 4.3 Morphological classification using non-parametric methods
here, Xi represents all pixel value , is the average of all Xi, is total number of pixels. All pixel values are arranged in an increasing order before calculating the . For a homogeneous system, converges to 0, whereas = 1 when the entire light of a galaxy is concentrated in a pixel (Glasser, 1962).
The parameter corresponds to the second-order moment of light for the brightest region in the galaxy (Lotz et al., 2004). The parameter is calculated by summing over the brightest pixel values () until the sum corresponds to 20% of the total galaxy flux. This value is normalized by Mtot corresponding to the second-order moment of the total light in a galaxy. These quantities are defined as:
here, xc, yc corresponds to galaxy central coordinates.
@@ 5 FUV emission in filament galaxies
The FUV SFRs would serve as an excellent tool to identify the most recent star-forming galaxy in our sample. As highlighted in the previous section, apart from UVIT-detected galaxies, we find a set of filament galaxies that do not emit in FUV above a threshold of 1.5. These galaxies are termed as FUV undetected. We perform fixed aperture photometry with an aperture radius of 1.5′′ on the FUV images at the centroid positions of the undetected galaxies. The magnitudes, thus obtained, were aperture corrected up to the saturation radius of a point source. The average aperture corrected magnitudes were approximately 1 - 1.5 mag higher than the FUV 3 limiting magnitude. These magnitudes are used in the subsequent analysis in this work.
@@
We calculated internal extinction corrected FUV SFRs for all (UVIT detected and undetected) filament galaxies using the following relation taken from Kennicutt & Evans (2012):
@@
where LFUVis the intrinsic FUV luminosity. We determine the value of internal extinction by the method discussed earlier in section 4.2. The ratio between stellar-to-gas dust attenuation is assumed to be 0.44 (Calzetti & Waller, 1997). Note that the stellar dust attenuation is affected by large-scale dust geometry in the case of galaxies hosting massive star-forming clumps. For such galaxies, the stellar-to-gas dust attenuation ratio may rise. Hence, a deviation from the local dust attenuation relation is expected. In other words, using the aforementioned local dust calibration relation, we may underestimate SFR from galaxies with the most clumpy FUV emission. The median FUV SFRs of UVIT detected (non-AGN) filament galaxies is 1.4 M⊙ yr-1, similar to the Milky Way galaxy. The stellar mass of filament galaxies was procured from GALEX-SDSS-WISE Legacy Catalog (GSWLC, Salim et al., 2016). Table 2 contains the FUV magnitudes, SFRs, stellar mass, and Dfil for all filament galaxies detected in UVIT.
5.1 SFRFUV SFR144MHz relation
As discussed above, the FUV emission is subjected to high dust attenuation due to the presence of dust clouds in the interstellar medium (ISM). The absorbed emission is re-emitted in Infrared (IR) wavebands ( between 8 m and 1000 m), hence, a combination of FUV and IR fluxes gives a precise estimate of the SFR in galaxies (Kennicutt & Evans, 2012; Buat et al., 2019). However, deep and well-resolved IR data is unavailable for our UVIT observed fields. The radio emission from star-forming galaxies is a promising alternative to estimate recent SFR. The emission at this wavelength in the star-forming galaxies is attributed to synchrotron radiation from accelerated electrons by the supernovae remnants, which is the end stage of short-lived massive stars and free-free emission from ionized gas around massive stars (Condon, 1992; Cram et al., 1998). Most importantly, this emission is not impacted by dust attenuation. We use low-frequency radio continuum observation (120 - 168 MHz) from LOw-Frequency ARray (LOFAR) Two-meter Sky Survey DR 2 (LoTSS; Shimwell et al., 2022) to estimate SFR144MHz for the FUV-detected filament galaxies. The survey has achieved a sensitivity upto 100 Jy and angular resolution of 6′′. The relationship between radio luminosity (L144MHz) and SFR at a short timescales ( 100 Myr) for a star-forming galaxy is well-established (Gürkan et al., 2018; Smith et al., 2021).
We detected nine filament galaxies in the LOFAR survey, of which three host AGN. The L144MHz (in W Hz-1) for our sample were calculated using total flux density (mJy) from LOFAR source catalog assuming a spectral slope (Hardcastle et al., 2016). The L is calibrated to SFR following Calistro Rivera et al. (2017). The distribution of SFRFUV and SFR for star-forming (circles) and AGN-host (crosses) filament galaxies is shown in the left panel of Figure 4. The color bar in the diagram denotes stellar mass. The non-AGN host galaxies with SFR M⊙ yr-1 from our sample were not identified in the LOFAR catalog. We highlight that both SFR indicators are in close agreement with each other for star-forming galaxies, while SFR144MHz in active galaxies is exorbitantly high compared to its FUV counterpart. The radio emission in such galaxies is powered by synchrotron emission from accreting central black holes.
5.2 sSFRFUV versus M⋆ as function of Dfil
The star-forming properties of galaxies depend on multiple factors, such as the halo-spin alignment of galaxies and Dfil (Kuutma et al., 2017; Mahajan et al., 2018; Ganeshaiah Veena et al., 2018, 2021). In this section, we discuss the variation in star-forming properties of galaxies with Dfil. The cross-sectional region of a filament is composed of three zones. The innermost ‘stream zone’ channelizes cold gas to massive galaxies, followed by the ‘vortex zone’ where the velocity is dominated by quadrupolar vortex structure risen due to the inflow of gas through sheets towards the center of filament, and the outermost ‘thermal zone’ where inflowing gas is decelerated due to thermal pressure (Lu et al., 2024). Studies investigating the connection between star formation quenching in the LSS have produced mixed results. While some reports suggest the cessation of star formation near the filamentary structures (Kuutma et al., 2017; Kraljic et al., 2018; Singh et al., 2020), a few observations report enhancement in star formation near filaments (Kleiner et al., 2017; Vulcani et al., 2019).
The right panel of Figure 4 shows FUV specific-SFR (log(sSFR) log(SFR) log(M⋆)) vs. M⋆ distribution for all filament galaxies targeted in our UVIT survey. We use aperture-corrected FUV magnitudes to calculate FUV SFRs in the case of UVIT undetected galaxies. The color bar indicates the Dfil for each galaxy. The AGN-hosts in our sample are marked with a maroon cross in the figure. These galaxies are removed from the analysis hereafter.
@@ 5.1 SFRFUV SFR144MHz relation
As seen in the figure, we find two distinct populations of galaxies in the filaments – star-forming and quiescent. A majority of the UVIT undetected galaxies lie closer to the filament spine. The non-detection of galaxies near the spine could be associated with the extra-galactic gas supply cutoff deep inside filaments suggested by Kuutma et al. (2017), which may lead to decline in recent star formation. We find that sSFRFUV and Dfil positively correlate for our combined sample. In other words, sSFR in galaxies increases with an increase in Dfil.
@@
Notably, most intermediate and high stellar mass galaxies detected in our survey are star-forming. It is well-established that several stellar mass quenching processes, namely gravitation quenching (Genzel et al., 2014), AGN feedback (Croton et al., 2006), active in massive galaxies lead to a decline in the ongoing star formation. A few studies on filaments establish that the inter-filamentary gas can replenish the HI content of filament galaxies (Kereš et al., 2005). Kleiner et al. (2017) reported the evidence of cold-mode gas accretion in several galaxies. This effect is more prominently observed in high-mass galaxies due to their deep gravitational potential well. The accumulated cold gas can further ignite the star formation activities in the massive filament galaxies. However, the likelihood of gas accretion in the sampled filament galaxies needs further investigation.
5.3 Comparison with galaxies from different LSSs
One of the key objectives of this work is to investigate whether the star-forming properties of filament galaxies are different from those existing in other LSSs, namely the voids and clusters. Most of the (non-AGN) galaxies detected in our FUV survey lie on the massive side of the stellar mass range (109.9 M M 1011. M⊙) while the median log(sSFRFUV) yr-1 for filament galaxies. For a small sample of void galaxies between stellar mass (1010 M⊙ to 1010.5 M⊙) taken from Void Galaxy Survey (VGS; Kreckel et al., 2012), the log(sSFR) yr-1, higher than our filament sample. Similarly, Conrado et al. (2024) report slightly elevated sSFRs in void galaxies compared to wall and filament galaxies.
@@ 5.3 Comparison with galaxies from different LSSs
Galaxy clusters are another LSS known to host the most massive galaxies in the Universe. We compare sSFR for filament and cluster galaxies in a fixed stellar mass bin (1010.0 M M 1011.0 M⊙). We use the RedMaPPer catalog (Rykoff et al., 2016) for identifying cluster galaxies in the nearby Universe (z 0.1). We select galaxies with cluster membership probability greater than 50%. The FUV magnitudes of cluster members were obtained from the GALEX AIS catalog. The galaxies with FUV magnitude error less than 0.35 mag are included in the sample. After applying all the cuts, we obtain a small sample of 18 (non-AGN) cluster galaxies. Using the FUV magnitudes, we calculate internal dust extinction FUV SFR using Equation 4. The value of internal reddening in the galaxies is deduced using the Balmar decrement ratio method (see Section 4.2). We create another dataset of cluster galaxies using deep GALEX observation of Coma cluster (Hammer et al., 2010). The observations reach a detection limit of 25 mag in FUV, equivalent to our UVIT observation. We identify cluster galaxies from GALEX observation following the structure of the Coma cluster defined in Colless & Dunn (1996). The FUV magnitudes were corrected for internal dust extinction using AFUV = 1.250.25 mag (Cortese et al., 2008) before SFR estimation. The stellar mass corresponding to both set of cluster galaxies were taken from GSWLC. We separate purely star-forming galaxies and AGNs using the BPT diagram from both sets and apply the stellar mass limit.
As shown in Figure 5, we find that filament galaxies have higher values of sSFR than both sets of cluster galaxies, implying that the filament galaxies are more star-forming than their counterparts in clusters. The observed results could arise from a slower gas depletion time scale in the former environment than the latter one, as pointed out by Castignani et al. (2022). Furthermore, the cold mode gas accretion from inter-filamentary channels or gas-rich mergers could rejuvenate the star formation in massive galaxies. It is worth noting that Parente et al. (2024) presented a similar trend for galaxies with M⋆ 1010.8 M⊙, showing higher sSFR with decreasing environment density; the star formation remains unaffected by the environment for more massive systems.
@@ 5.4 FUV morphology of filament galaxies
The FUV morphology of galaxies is an efficient probe to identify the clumpy star-forming regions inside them. We utilize our well-resolved UVIT observation to investigate the FUV morphology of filament galaxies and compare them with their optical counterparts. Figure 6 illustrates FUV images of all UVIT-detected filament galaxies. The blue contours on the figure show the SDSS-r band profile of galaxies. The three contour levels in blue represent 10, 5, and 1.5 level, where represents the mean standard deviation in the SDSS r-band observation. The surface brightness of the outermost contour is 24.14 mag arcsec-2. The galaxies are serially assigned a number () according to Table 2.
@@
We show the FUV and optical centroids of the galaxies using cyan and red crosses, respectively. The young stellar population inside galaxies represented by FUV emission mostly aligns with the main stellar population traced by optical emission. On the contrary, a similar analysis done for cluster galaxies by Mahajan et al. (2022) highlights that the FUV morphology for a sample of cluster galaxies was skewed relative to their optical emission primarily due to ram-pressure stripping (Gunn & Gott, 1972) rampantly observed in the clusters. The result establishes that filament galaxies may not witness extreme environmental disturbances, which are prevalent in clusters. However, the FUV emission appears clumpy in the filament galaxies and shows considerable size variation. For a few galaxies, the emission is asymmetric within the galaxy, e.g., . Moreover, we observe FUV emission in some galaxies beyond the outermost optical contour, e.g., . These observations hint at the possibility of recent gas accretion episodes. A thorough comparison between deep optical and UV observation of the filament galaxies is required for confirming the existence of extended-UV disks (Thilker et al., 2007; Borgohain et al., 2022).
@@
As discussed in the previous section, massive galaxies in filament exhibit slower star formation quenching than those in the clusters. The FUV disk sizes of most of the filament galaxies in the sample are slightly extended or comparable to their optical emission. These galaxies are mostly detected in the outer edge of filaments (D Mpc ). Previous studies establish that massive and older galaxies reside closer to the filament center (Chen et al., 2017; Kuutma et al., 2017). In our survey, massive galaxies near the filament axis were undetected or detected with weak UV emission (see in Figure 6; also see upper left panel of Figure 3). An earlier study analyzing the HI gas content of galaxies in the filament environment reported that at fixed stellar mass and color, the galaxies closer to the filament spine are increasingly HI deficient (Odekon et al., 2018). This finding corroborates our result that the galaxies closer to the spine do not show a signature of recent star formation.